OpenStudy (amenah8):

Will fan and Medal!! I just need help on problems 41, 49, 52, and 58. I've done all the rest! Click for link to problems

OpenStudy (amenah8):

Please, even if you can only help with one! I can always just post another question for the rest

OpenStudy (peachpi):

41. The book hits the ground at p(t) = 0. Solve for t. Also, v(t) = p'(t). Use this to find the speed when p(t) = 0.

OpenStudy (peachpi):

49. Seems like part A is asking you to pick points off the graph and use them to find 2 estimates of the area under the curve from 0 to 4. (left, right, midpoint, trapezoid, etc) B. \[\int\limits_{0}^{4}a(c)~dc\]

OpenStudy (peachpi):

52. If you assume forward is the positive direction, then the car moves forward when v(t) > 0, backward when v(t) < 0 and is not moving when v(t) = 0.

OpenStudy (peachpi):

58. that integral and R(T) tell you the same thing, the barrels of oil pumped in T days. So R(5) gives how many barrels were pumped in 5 days.

OpenStudy (amenah8):

ok i understand all of your explanations except for 41.

OpenStudy (amenah8):

@peachpi am I solving for t when p(t)=0 or am I solving for t for the equation p(t)=1250-16t^2?

OpenStudy (peachpi):

you're setting p(t) equal to 0 then solving for t p(t) = 1250 - 16t^2 p(t) = 0 1250 - 16t^2 = 0

OpenStudy (amenah8):

@peachpi x=8.84??

OpenStudy (peachpi):

yes

OpenStudy (peachpi):

that's the time when it hits the ground. Now integrate to find the speed at this time

OpenStudy (amenah8):

find the derivative of 8.84? or the equation?

OpenStudy (peachpi):

the equation

OpenStudy (amenah8):

do i plug in 8.84 for t?

OpenStudy (peachpi):

yes

OpenStudy (peachpi):

and that should give you velocity. I think you can just make velocity positive to get speed

OpenStudy (amenah8):

-.3296

OpenStudy (amenah8):

and then make it positive: .33

OpenStudy (amenah8):

so velocity is about .33m/s?

OpenStudy (amenah8):

and that would be the exact speed for when the book hits the ground?

OpenStudy (peachpi):

way too low. what was your derivative?

OpenStudy (peachpi):

should have had v(t) = p'(t) = -32t. then plug in 8.84 for t

OpenStudy (amenah8):

OH woops I plugged in 8.84 for 1250-16(8.84)

OpenStudy (amenah8):

ok i got what you did

OpenStudy (amenah8):

and make it positive?

OpenStudy (amenah8):

so v(t)=282.88m/s??

OpenStudy (peachpi):

yes

OpenStudy (amenah8):

hooraaay! what about for a, where i need 3 approximations for speed using different size intervals? is that left and right hand sums or something else?

OpenStudy (peachpi):

I think they just want you to find the average rate of change on 3 different intervals. The smaller will be the most accurate because it's a smaller snapshot and should be closer to the instantaneous speed you just found. Since you're investigating what happens at t = 8.84 I'd use these intervals 7 s to 8.84 s 8 s to 8.84 s 8.5 s to 8.84 s

OpenStudy (amenah8):

ok so for example on the interval from 7 to 8.84 seconds just plug both of those numbers into the equation? like over and under estimates? and the most accurate one would be the third interval

OpenStudy (amenah8):

but i don't need to take the derivative correct? or do i in order to get velocity?

OpenStudy (peachpi):

no you don't need to take the derivative. Avg rate of change will give you an estimate for the velocity. Basically you use this equation \[aroc=\frac{f(b)-f(a)}{b-a}\]

OpenStudy (amenah8):

so 8.84-7 divided by what?

OpenStudy (peachpi):

For the first interval p(7) = 466 p(8.84) = 0 \[\frac{p(8.84)-p(7)}{8.84-7}=\frac{0-466}{8.84-7}=-253.3\] Then make 253.3 for speed

OpenStudy (amenah8):

so the speed for that interval is 253.3m/s?

OpenStudy (peachpi):

yes

OpenStudy (peachpi):

you got it from here?

OpenStudy (amenah8):

yes i think so ! thank you soooo much

OpenStudy (peachpi):

you're welcome